# BOUQuET [Uncaptioned image]: dataset, Benchmark and Open initiative for Universal Quality Evaluation in Translation

### v1

@@ 2.3 Languages
ind Latn Indonesian Austronesian Malayo-Polynesian
@@ 3 Benchmark
Statistics
In total, BOUQuET currently contains 1,750 sentences. These sentences are split by doing a stratified selection among source languages and domains into development and test sets with an equivalent number of sentences. A portion of the development set (522) is intended to initially be kept hidden. Results in the following sections are presented with the test split of 889 sentences.
Table 3 provides a comparison of several relevant statistics from BOUQuET and the closest related datasets including FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (NLLB-Team et al., 2022).
Dataset Sent. Parag. Reg. Dom. Lang. Dyn.
FLORES-200 3,003 () $\times$× Wikipedia, News, Travel guides 200 $\times$×
NTREX-128 1,997 $\times$× News 128 $\times$×
NLLB-MD 9,000 $\times$× $\times$× Chat, News, Health 6 $\times$×
BOUQuET 1,750 Fiction, Conversation, Social media posts/comments 8+
Tutorials, Website, Reflection pieces, Miscellaneous
Table 3: Main statistics from MT evaluation datasets including BOUQuET: number of sentences, paragraph-information, registers, domains, languages, dynamism.
@@
The performance of the model in a new or unseen dataset depends on the similarity between the dataset that was used to fit the model and the new dataset. In order to more broadly evaluate an MT model, we are interested in covering a different set of domains. We compare the domain coverage of BOUQuET with that of FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (Stolte et al., 2024) . To do this comparison, we represent data (multi-domains public datasets from Figure 2 in grey, each alternative dataset FLORES-200, NTREX-128 and NLLB-MD and BOUQuET) with SONAR (Duquenne et al., 2023). From SONAR vectors, we do a PCA-dimensionality reduction, see Figure 3. All domains from Figure 2 are represented in grey. Then, Figure 3 from left to right compares BOUQuET against FLORES-200, NTREX-128 and NLLB-MD, respectively. We qualitatively observe that BOUQuET covers a wider range of domains.
Additionally, to quantify this coverage, we measure the overlap between each dataset with each of the domains using two metrics, the Wasserstein distance (implemented with POT library [^5] https://pythonot.github.io/) and the Silhouette coefficient (implemented with sklearn [^6] https://scikit-learn.org/stable/). The Wasserstein Distance (WD), also known as the Earth Mover’s Distance (EMD), is a metric that measures the "effort" required to transform one probability distribution into another. Lower results indicate a higher similarity between clusters. The lowest consistent results for all domains are obtained with the data set BOUQuET. Alternatively, the Silhouette Score ($S$)  [^7] https://en.wikipedia.org/wiki/Silhouette_(clustering) measures how well each data point fits within its assigned cluster compared to other clusters, it ranges between -1 and 1 and the higher the coefficient, the more coherent the clusters are. Again, BOUQuET shows a lower score, consistent across domains, meaning closer representation across domains.
Figure 3: Domain representation and overlap with datasets (FLORES-200, NTREX-128, NLLB-MD)
Narration Website Opinion Dialog Social media Social media How-to Misc.
comments posts
WD SC WD SC WD SC WD SC WD SC WD SC WD SC WD SC
FLORES-200 0.32 0.05 0.32 0.03 0.32 0.02 0.31 0.11 0.32 0.09 0.31 0.08 0.31 0.06 0.33 0.03
NTREX-128 0.32 0.04 0.31 0.03 0.32 0.01 0.31 0.09 0.31 0.07 0.31 0.6 0.31 0.06 0.33 0.03
NLLB-MD 0.31 0.04 0.31 0.02 0.31 0.00 0.30 0.08 0.31 0.06 0.30 0.05 0.30 0.05 0.32 0.02
Bouquet 0.29 -0.02 0.29 -0.02 0.30 -0.03 0.28 0.02 0.29 0.01 0.28 0.00 0.28 0.00 0.30 -0.01
@@
Table 4: Wasserstein Distance (WD) and Silhouette Score ($S$) for each domain and dataset. Lower WD and $S$ indicate better representation of the domain. Best results in bold.
MT Benchmarking
To help the reader understand why the dataset is useful, we present preliminary results to demonstrate its use for its intended purpose: MT benchmarking. We evaluated on translation systems (GPT-4o (Wu et al., 2024), Llama-3.1 (Llama-3.3-8B)(Touvron et al., 2023), and NLLB (NLLB-3.3B) (NLLB-Team et al., 2022)) using two metrics: Comet (CometKiwi-da-xl, range 0-1 and $\uparrow$ better) (Chimoto and Bassett, 2022) and MetricX (MetricX-24, range 0-25 and $\downarrow$ better) (Juraska et al., 2024), following WMT 2024 (Kocmi et al., 2024). See results in Table 5. We observe that BOUQuET scores consistently higher than NLLB-MD on average and for all models, suggesting that BOUQuET is easier to translate. This is an advantage for the open initiative, since the complexity of current MT test sets makes it harder to ask the community to participate in translations as it requires a high-level of expertise. Rankings across models nor languages is not preserved, which hints that BOUQuET may be posing different challenges to the models. We need to further investigate which linguistic challenges BOUQuET is adding. On this ranking, if we look at average BOUQuET, LlaMa is the worst system according to Comet but the best system according to Metricx. This does not happen in NLLB-MD, where the ranking is consistent across metrics. If we look per language, ranking across metrics is consistent for 4 out of 7 languages for NLLB-MD and for 1 out of 7 languages for BOUQuET. This raises a concern and may suggest that MT metrics themselves do not generalize well to the broader domains/registers in BOUQuET. However, this requires further investigation and we want to specifically explore different alternatives to use BOUQuET such as exploring the performance by domain and register. Note that we present only preliminary experiments, see Section 5 for more details on ongoing experiments.
M cmn deu fra hin ind rus spa AVG
COM($\uparrow$) MetX($\downarrow$) COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX
NLLB-MD
GPT4o 0.76 14.34 0.77 6.80 0.76 13.86 0.68 6.63 0.79 16.16 0.78 3.96 0.80 13.99 0.76 10.81
Llama 0.71 14.38 0.71 6.96 0.72 14.00 0.57 7.07 0.75 16.21 0.71 4.00 0.75 14.11 0.70 10.96
NLLB 0.61 14.80 0.70 7.01 0.70 13.98 0.64 6.72 0.73 16.20 0.72 4.43 0.76 14.06 0.69 11.03
BOUQuET
GPT4o 0.79 1.99 0.80 2.79 0.78 5.57 0.76 3.37 0.84 4.07 0.81 3.60 0.83 4.74 0.80 3.73
Llama 0.75 2.06 0.76 2.07 0.74 4.68 0.72 3.19 0.83 3.54 0.75 3.26 0.80 4.07 0.76 3.27
NLLB 0.67 2.19 0.79 2.44 0.77 5.30 0.74 3.49 0.82 3.94 0.78 3.40 0.67 4.42 0.77 3.60
Table 5: Results on Quality Estimation from English to 7 Source-BOUQuET languages on BOUQuET evaluation dataset .
4 Open initiative
@@
Source-BOUQuET is intended to be translated into any written language. This ambition can only be achieved with the support of the community. For this, we have organized an open collaborative effort which involves language communities that are interested in contributing to this effort.
The purpose of this open initiative is to collect translations from Source-BOUQuET. To collect these annotations, we have chosen the ARGILLA tool [^8] https://argilla.io/ which is a collaboration platform for building datasets, designed for use by AI engineers and domain experts.
Together with setting the Source-BOUQuET in this tool, we have designed annotation guidelines which very much resemble those from FLORES-200 (NLLB-Team et al., 2022).
This open initiative is available in HuggingFace https://huggingface.co/spaces/facebook/bouquet and it is starting as part of the Meta Language Technology Partner Program [^9] https://docs.google.com/forms/d/e/1FAIpQLSdzcRdtkQCuTrXw727DgJgWbOPKDj5v0bArgGfQUTT6sEopFw/viewform publicly launched in AI Summit [^10] https://www.elysee.fr/en/sommet-pour-l-action-sur-l-ia.
@@ 5 Conclusions and Next Steps
In this paper, we have presented the Source-BOUQuET dataset and the attached open initiative. This dataset is available in 8 languages. We have shown that consistent BOUQuET gains in domain diversity by in two different metrics while keeping complexity lower than its competitors. The latter is specially relevant to simplify the translation for non-experts that may join the open initiative. This paper and Source-BOUQuET is only a fraction of the ambitious project that we are starting by launching this open call for community efforts. Please join us in making Universal Quality Evaluation in Translation available in any language!
Limitations
BOUQuET is work in progress; the BOUQuET dataset is still limited in number of languages and translations. The benchmarking is also limited in several axes (dataset comparison, models and metrics evaluated). However, we prioritized starting the BOUQuET open initiative in coordination with the Meta Language Technology Partner Program launch so that we could start raising the collaboration interest. As OmnilingualMT team, we continue actively investing on it, see next steps that we are planning to release soon.
Coming Soon
BOUQuET is actively evolving and we are currently working on making the following progress:
  • Adding 250 sentences in Source-BOUQuET in Egyptian Arabic.
  • Complete Source-BOUQuET in all 7 languages from Table 2 plus Egyptian Arabic.
  • The creation guidelines plus translation guidelines in all 8 pivot languages.
  • Designing quality control for each of the contributions and adding new languages to the incremental releases of BOUQuET.
  • Extending the Benchmarking by further showing the capabilities of BOUQuET, e.g. increasing the evaluation of linguistic signals over its alternatives (FLORES-200, NTREX-128, and NLLB-MD). For this, we propose to explore the performance in different domains and see if we get difference in the performance ranking of translation systems.
Acknowledgments
We would like to extend our sincere gratitude to Luis Cadavid, Radhi Datla, Suci Fitriany, Christine Haugen, Stephanie Mauro, Carissa Ikka Pardamean, Gauri Shanware, and Jyoti Singh for their invaluable contributions to the work presented in this paper.
@@ A.1 Overview
The Bouquet-source dataset needs to include 250 sentences in each of its source languages. This means that each linguist needs to create (i.e., create from scratch, not copy; see Section 2.4 above) 250 original sentences. These sentences must be:
@@ A.2 Paragraphs
The linguist will receive a template in the form of a spreadsheet, in which paragraph structures are designed and laid out. The template specifies the exact number of paragraphs and the exact number of sentences for each of the paragraphs. Each paragraph is given a unique paragraph ID (e.g., P01, P02, P15). Each sentence within each paragraph is also given a serial, non-unique ID (e.g., S1, S2, S3). The existing structure must not be modified in any way.
@@ A.3 Domains
The template is divided into 8 domains:
@@
You must produce the set number of sentences for each of the domains; the domain / paragraph / sentence structure of the template cannot be changed.
@@ A.4 Language Register Information
When creating sentences, make sure that the register of language being used is representative of the most expected and appropriate register for the situation. We understand that several registers could be possible; please use discretion when selecting a register, while making sure that the chosen register is among the most expected and appropriate. To help you make a determination, we define 3 main functional areas of language register:
@@ A.6 Violating Content
While creating sentences, please be sure to avoid inserting violating content. Violating content is language that can fall under one (or more) of the below categories:
@@ A.7 Step-by-Step Description of Tasks
Please use the provided template to follow this step-by-step description.
@@
Column U: Creator_Translator-ID Please insert your ID here, if it isn’t prepopulated.
@@ A.8 Additional Guidance on Domain-Specific Content
Dialogues, especially those inserted in long creative writing (such as novels), often include the name of the speaker or a cue mark (e.g., — ), and sometimes quotation marks. When creating sentences for conversations, please feel free to invent names for speakers or to label speaker turns (e.g., A, B). Please place the names or speaker reference in markup tags, similarly to this: <Name:>or <A:>.
@@
Emojis: There are emojis frequently in some social media and messaging domains. There should therefore be a representation, albeit very limited but the point isn’t to know how the system translates emojis because there are no real agreed ways to translate them across hundreds of languages.
Social media comments: Okay if the structure is flattened. Okay to have some tags but not absolutely necessary and no need to insert them everywhere.

### v2

@@ 3.4 Multi-way extension to Source-BOUQuET languages
ind Latn Indonesian Austronesian Malayic
@@
3.5 Overall Statistics
In total, BOUQuET contains 2,000 sentences. These sentences are split by making a stratified paragraph-level selection among source languages and domains into development, test and evaluation sets. Initially, the evaluation set (632/144 sentences/paragraphs) is intended to be kept hidden. Figure 2 shows the representation of registers (top) and domains (bottom) in the non-hidden splits. Labels for each of the combinations of register options are created by concatenating the lowercase letters used as unique identifiers (see details of these register options in the Appendix B). For example, a register characterized as impersonal (in connectedness), composed (in preparedness), and equal-assumed (in social differential) is labeled ica.
Figure 2: Registers (top) and domains (bottom) representations in development and test partitions.
The results in the following section are presented with the test split of 864/200 sentences/paragraphs.
@@ 4 Benchmark
We benchmark BOUQuET in two dimensions: domain representation and machine translation. The former quantifies how representative BOUQuET is of public datasets of multiple domains compared to other evaluation datasets. The latter addresses how several MT systems are ranked with BOUQuET compared to other evaluation datasets.
@@
The performance of the model in a new or unseen dataset depends on the similarity between the dataset that was used to fit the model and the new dataset. We compare the domain coverage of BOUQuET with that of FLORES+, NTREX-128 and NLLB-MD. To do this comparison, we take a random sample of 2,000 sentences (which seems to be a sufficiently large sample of the embedding space for score stability) from each of the domain datasets from Figure 1; as well as 2000 from each alternative dataset FLORES+, NTREX-128, NLLB-MD,and BOUQuET. We create vector representations of each sentence in previous datasets with SONAR (Duquenne et al., 2023). From SONAR vectors, we do a PCA-dimensionality reduction, fitted upon the combined multi-domain set, see Figure 6 (Appendix F). Public domains from Figure 1 are represented in grey; alternatives evaluation datasets are represented in blue and BOUQuET is represented in red. Figure 6, from top to down, compares BOUQuET against FLORES-200, NTREX-128, NLLB-MD, respectively. We qualitatively observe that BOUQuET covers a wider range of domains. Additionally, to quantify this coverage, we measure the overlap between each dataset with each of the domains using the Wasserstein distance (implemented with the POT library [^7] https://pythonot.github.io/). The Wasserstein Distance (WD), also known as the Earth Mover’s Distance (EMD), is a metric that measures the ”effort” required to transform one probability distribution into another. Lower results indicate a higher similarity between clusters. This distance is run on the full 1024-dimensional SONAR embedding vectors, without applying any kind of dimensionality reduction (PCA). Some domain sets and evaluation sets were several orders of magnitude larger than each other. Sampling all down to the same size (2,000) makes the metric computable in a reasonable amount of time and removes any sensitivity to class imbalance in the distribution distance metric. Figure 3 shows that the lowest consistent results are obtained for all domains with BOUQuET.
@@
Figure 3: Wasserstein Distance (WD) for each domain and dataset. Lower WD indicate better representation of the domain.
Machine Translation
To help the reader understand why the dataset is useful, we present preliminary results to demonstrate its use for its intended purpose: MT benchmarking. We evaluate 14 translation systems: Llama-3 (Llama3.1-8B, Llama3.2-3B, Llama3.3-70B)(Touvron et al., 2023), Tower (TowerInstruct-7B-v0.2) (Rei et al., 2024), Aya (Aya101-13B, Aya-Expanse-8B) (Dang et al., 2024), Babel (Babel-9B-Chat) (Zhao et al., 2025), Cohere (CohereLabs-command-r7b-12-2024), Eurollm (EuroLLM-9B-Instruct) (Martins et al., 2024), MADLAD (MADLAD-3B-MT and MADLAD-10B-MT) (Kudugunta et al., 2023), Mistral (Mistral-7B-Instruct-v0.3) [^8] https://docs.mistral.ai/getting-started/models/models_overview/, Qwen (Qwen2.5-7B-Instruct) (Bai et al., 2023) and NLLB (NLLB-3.3B) (NLLBTeam, 2024). We select the models as ones with open weights, focusing primarily on moderate sizes (about 10B) and variety of architectures. Following the official evaluation metrics of WMT 2024 (Kocmi et al., 2024), we use two automatic metrics: CometKiwi (CometKiwi-da-xl, range 0-1 and $\uparrow$ better, COM) (Chimoto and Bassett, 2022) and MetricX (MetricX-24-hybrid-xl-v2p6, range 0-25 and $\downarrow$ better, MetX) (Juraska et al., 2024). We include in the benchmarking datasets that cover Source-BOUQuET languages (FLORES+ and NTREX-128).
Table 4 shows that BOUQuET scores consistently higher than other datasets on average, suggesting that BOUQuET is easier to translate. This is an advantage for the open initiative, since the complexity of current MT test sets makes it harder to ask the community to participate in translations as it requires a high-level of expertise.
Rankings across models and datasets is not preserved, which hints that all datasets may be posing different challenges to the models. Rankings is computed as counting when a system is similar in the same position according to CometKiwi. This ranking and Pearson correlation on the CometKiwi is dissimilar for datasets evaluated at the sentence-level, with BOUQuET being the most different. This difference is enlarged when evaluating at the paragraph-level where number of swaps increases and, coherently, Pearson correlation decreases, meaning that datasets pose different challenges to models. We need to further investigate which linguistic challenges BOUQuET is adding. However, best two systems are consistent across datasets and level of evaluation (sentence and pargraphs) being those the largest model (Llama3.3-70B) and Aya-e-8B.
NLLB-3.3B has a higher variation between being evaluated at the sentence or paragraph-level, which makes sense since it is the only one trained with sentence-level data.
Model Bouquet Flores Ntrex BouquetP NtrexP
COM MetX COM MetX COM MetX COM MetX COM MetX
NLLB-3B 0.68 2.1 0.66 2.56 0.65 2.97 0.59 3.71 0.29 14.1
aya101-13B 0.67 2.02 0.63 2.65 0.63 3.14 0.58 3.29 0.24 13.17
aya-e-8B 0.69 1.75 0.65 2.9 0.67 2.45 0.64 2.42 0.34 8.7
babel-9B 0.67 2.33 0.65 2.66 0.63 3.36 0.61 3.4 0.32 10.39
cohere-7B 0.67 2.15 0.65 2.89 0.64 3.01 0.61 3.2 0.32 9.61
eurollm-9B 0.67 2.33 0.65 2.89 0.61 3.64 0.61 3.64 0.31 10.08
madlad-10B 0.63 2.74 0.64 2.72 0.63 3.35 0.41 6.76 0.15 15.99
madlad-3B 0.63 2.85 0.63 2.94 0.61 3.67 0.37 6.71 0.49 5.29
mistral-7B 0.54 4.29 0.51 5.69 0.49 6.12 0.49 6.64 0.24 10.96
qwen-7B 0.59 3.25 0.6 3.75 0.59 4.21 0.57 4.5 0.52 4.93
Llama3.1-8B 0.66 2.36 0.64 2.82 0.63 3.27 0.6 3.33 0.32 10.17
Llama3.2-3B 0.59 3.59 0.57 4.34 0.55 4.89 0.52 5.52 0.27 12.67
Llama3.3-70B 0.7 1.85 0.68 2.21 0.67 2.59 0.63 2.72 0.35 9.76
Tower-7B 0.58 3.69 0.56 4.19 0.56 4.35 0.49 5.65 0.28 12.22
Bouquet-Flores Flores-Ntrex Ntrex-Bouquet BouquetP-NtrexP
Swaps 3 4 7 11
Pearson Cor. 0.95 0.99 0.95 0.92
Table 4: Averaged Results $X$-to-$X$ 9 Source-BOUQuET (8 pivot plus English) languages for BOUQuET, FLORES+, NTREX-128 at the level of sentence 2 columns on the left and at the level of paragraph 3 columns on the right. Number of ranking swaps (a system not being in the same position according to CometKiwi) from each dataset compared to the other two (in similar sentence or paragraph-level) and Pearson correlation indicate that while datasets report similar results at sentence-level, being BOUQuET the most different, it is not the case for paragraph-level where the ranking of systems varies by a larger amount.
Figure 4 shows results of the 3 best systems averaged across language directions, evaluated at the sentence-level, per domains. Worse performing domains are comments, conversations, how-to and narration. Best performing domains are web and other miscellaneous, reflection and social posts. Appendix F reports more detailed results on BOUQuET per language and domains.
Figure 4: Best performing models and their results in each of the BOUQuET domains .
5 Beyond commissioning translations: Open initiative
Source-BOUQuET is intended to be translated into any written language. For this, we have commissioned an initial set of priority languages covering a variety of high and low-resource languages representing different geographical regions, linguistic families and scripts. See the list of languages currently covered by BOUQuET in Appendix D.
@@ 5 Beyond commissioning translations: Open initiative
However, it would be challenging to achieve our language coverage target to any language. This ambition can only be achieved with the support of the community. For this, we have organized an open collaborative effort which involves language communities that are interested in contributing to this effort.
The purpose of this open initiative is to collect translations from Source-BOUQuET. To collect these annotations, we have set a tool to collect annotations. Together with setting Source-BOUQuET in this tool, we use the annotation guidelines from Section 3.4 which very much resemble those from FLORES-200 (NLLBTeam, 2024) and which are available in the 9 BOUQuET languages. One of the advantages is that annotators can choose the source language from among one of the Source-BOUQuET languages. These languages have been chosen to cover a wide range of speakers, facilitating the task of annotation instead of depending on English bilingual speakers. This open initiative is available in https://bouquet.metademolab.com/.
@@ 6 Conclusions and Next Steps
In this paper, we have presented the Source-BOUQuET dataset and the attached open initiative. We have shown consistent gains in domain diversity in two different metrics while keeping complexity lower than its competitors. The latter is particularly relevant to simplify the translation for non-experts that may join the open initiative. We also provide MT results for the 8 languages in which Source-BOUQuET has been created. Although BOUQuET is currently totally completed for 55 languages (see list 6), this number is only a fraction of the language coverage ambition that we are pursuing by launching the open initiative for community efforts. Please join us in making Universal Quality Evaluation in Translation available in any language.
Beyond increasing in number of languages, BOUQuET is actively evolving, and we are currently working on designing quality control for each of the contributions and adding new languages to the incremental releases of BOUQuET and extending the benchmarking by further showing the capabilities of BOUQuET, e.g. increasing the evaluation of linguistic signals over its alternatives.
Limitations and Ethical Considerations
The BOUQuET dataset is still limited in the number of languages and translations.The benchmarking is quite complete (4 datasets comparison, 14 models and 2 metrics) but it can also be extended in several axes (linguistic analysis). However, the entire purpose of this work is to describe the dataset and open-initiative, while providing a minimal benchmarking. Authors expect the community to extend the benchmarking by further using this dataset for further exploration. Creators and commissioned translation’s annotators are paid a fair rate.
@@ A.1 Overview
The Bouquet-source dataset comprises 250 unique sentences in each of its source languages. This means that each linguist created (i.e., wrote from scratch, did not copy; see Section 2.4 above) 250 original sentences. These sentences were requested to be:
@@ A.2 Paragraphs
The linguist received a template in the form of a spreadsheet, in which paragraph structures were designed and laid out. The template specified the exact number of paragraphs and the exact number of sentences for each of the paragraphs. Each paragraph was given a unique paragraph ID (e.g., P01, P02, P15). Each sentence within each paragraph was also given a serial, non-unique ID (e.g., S1, S2, S3).
@@ A.3 Domains
The template was divided into 8 domains:
@@
The creators had to produce the set number of sentences for each of the domains; the structure of the template (domain / paragraph / sentence) could not be changed.
@@ A.4 Language Register Information
When creating sentences, the creators had to make sure that the register of language being used was representative of the most expected and appropriate register for the situation. When several registers were possible, the creators were asked to use discretion when selecting a register, while making sure that the chosen register was among the most expected and appropriate. To help them make a determination, we defined 3 main functional areas of language register:
@@ A.6 Violating Content
While creating sentences, the creators were asked to avoid inserting violating content. Violating content is language that can fall under one (or more) of the below categories:
@@ A.7 Step-by-Step Description of Tasks
Please refer to Table 5 for the step-by-step description of the tasks.
@@
Column U: Creator_Translator-ID Please insert your ID here, if it isn’t pre-populated.
Table 5: Step-by-step guidance.
@@ A.8 Additional Guidance on Domain-Specific Content
Dialogues, especially those inserted in long creative writing (such as novels), often include the name of the speaker or a cue mark (e.g., — ), and sometimes quotation marks. When creating sentences for conversations, the creators were let free to invent names for speakers or to label speaker turns (e.g., A, B); they were also asked to place the names or speaker reference in markup tags, similarly to this: <Name:>or <A:>.
@@
Emojis: As there are emojis frequently in some social media and messaging domains, some representation was also expected from the creators. However, the creators were asked to keep this representation very limited, as there are no real agreed ways to translate them across hundreds of languages.
Social media comments: The creators were told that they could keep the structure of those comments flat, and that including tags was not absolutely necessary, though it was permitted (even expected).